# Risk-Sensitive Agent Compositions

### v2

@@ Appendix A Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
This implies that $q\geq\text{VaR}_{\alpha}[L_{p}(Z_{p})]$. However, in the algorithm, we estimate the $(1-\alpha_{i})$-quantile for each edge empirically using $n$ samples. By the Dvoretzky-Kiefer-Wolfowitz (DKW) inequality [8, 23], with probability at least $1-\delta^{\prime}$, the empirical CDF $\hat{F}_{i}$ for each edge satisfies:
@@
Appendix B Benchmark environments and experimental evaluation setups
B.1 16-Rooms
@@ Appendix B Benchmark environments and experimental evaluation setups
This benchmark environment is borrowed from Jothimurugan et al. [18]. The state and action spaces are continuous and involves controlling a point mass to navigate between rooms while avoiding obstacles. A map of the environment is shown is shown in Figure˜5. Jothimurugan et al. [18] specify objectives in the form of task graphs where each vertex corresponds to a set of subgoal states and each edge represents a reach-avoid subtask of reaching the target vertex’s subgoal states. The task graph that we consider corresponds to the specification $\phi_{5}$ from Jothimurugan et al. [18] and is also visualized in Figure˜5. It consists of 13 vertices arranged in a sequence of 4 diamonds with 4 vertices in each diamond like in Figure˜1. The first diamond specifies starting from the light blue square and reaching the room marked with the dark blue square in Figure˜5 using either the top path while avoiding the red obstacle or using the bottom path. Similarly, the next diamond specifies the two paths to reach the green square, followed by the pink, and brown squares. Each edge in the task graph is also associated with a real-valued reward function that evaluates to at least zero when a trajectory of the controller satisfies the reach-avoid subtask. This is function of the distance between a trajectory and the obstacles and between the trajectory and the subgoal region. So if the reward is positive then it means that the trajectory stays at least a certain distance away from the obstacles and reaches at most a certain distance to the goal.
Mapping a task graph to an agent graph is simple: to define the loss functions along each edge, we simply consider the negative of the reward function so that an upper bound on the loss corresponds to a lower bound on the reward. The agents along each edge are RL policies trained to complete the reach avoid subtasks. The policies are trained using the ARS algorithm [22] using the same recipe as described in Jothimurugan et al. [18]. Since the training of policies are dependent on the initial state distribution, we train the policies lazily when required by the BucketedVaR algorithm by approximating the initial state distribution from samples from the previous policy.
@@
In Table˜1, since the estimated VaR at all levels for the 16-Rooms benchmark is negative, it implies that with probability at least $1-\alpha$, the reward along all the edges of the path are at least the positive estimated VaR, where $\alpha$ is the risk level.
B.2 Fetch
The Fetch robotic arm environment [26] from Gymnasium Robotics is a simulated robot arm environment with a gripper attached. The reach-avoid task graph is again taken from Jothimurugan et al. [18] and has 8 vertices with two paths. It corresponds to the PickAndPlaceChoice specification and involves moving the gripper close to an object, gripping it, picking it up, and moving it along one of two trajectories. The policies are trained using TD3 [10] with the hyperparameters specified by Jothimurugan et al. [18].
B.3 DroneNav
@@
B.4 BoxRelay
Figure 6: First-person view when carrying box in the BoxRelay environment which also serve as observations.
The environment and the agent graph details are presented in Section˜4. The policies are trained using the PPO [27] CNN-Policy algorithm from StableBaselines3 package with the default hyperparameters for 500,000 iterations for each edge policy. The initial state distribution was again estimated by taking rollouts of the previous edge policy. The loss functions in the agent graph are defined as follows where the vertices are numbered as shown in Figure˜2.
@@ B.4 BoxRelay
B.5 Experiment computational requirements
@@
Appendix C Experimental results
Table 1: Comparison of BucketedVaR with the optimal baseline algorithm on $10^{4}$ samples. We prefer empirical quantiles that are close to $1-\alpha$ where $\alpha$ is the risk level. 95% Clopper-Pearson CIs are specified for the empirical quantiles computed on a fresh set of $10^{4}$ samples.
@@ Appendix C Experimental results
Complete evaluation results of the BucketedVaR algorithm along with comparison with the baseline optimal algorithm are given in Table˜1.

### v3

@@ Appendix B Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
This implies that $q\geq\text{VaR}_{\alpha}[L_{p}(Z_{p})]$. However, in the algorithm, we estimate the $(1-\alpha_{i})$-quantile for each edge empirically using $n$ samples. By the Dvoretzky-Kiefer-Wolfowitz (DKW) inequality [Dvoretzky et al., 1956, Massart, 1990], with probability at least $1-\delta^{\prime}$, the empirical CDF $\hat{F}_{i}$ for each edge satisfies:
@@
Appendix C Benchmark environments and experimental evaluation setups
C.1 Maximizing safety during navigation and manipulation
We test our algorithm on the 16-Rooms and Fetch environments from Jothimurugan et al. [2021] who present a compositional RL framework to complete long-horizon objectives in continuous environments. An objective is specified as a task graph which is the same as an agent graph. In the task graph, every edge represents a reach-avoid subtask: reaching a target region while avoiding a set of dangerous states. They design an algorithm that finds the path that maximizes the probability of completing all the reach-avoid subtasks along the path where the control policies along the edges are trained using RL. Each reach-avoid subtask specifies a real-valued reward function that takes in a trajectory of execution of the policy and evaluates to at least zero if the trajectory successfully completes the subtask. Intuitively, the reward captures a distance metric between trajectories and reach-avoid sets. The loss function in the agent graph is defined as the negative of the reward function.
The 16-Rooms environment consists of 16 rooms arranged in a $4\times 4$ grid with doors between adjacent rooms. RL policies are trained to control a point mass to complete reach-avoid tasks. The agent graph has 13 vertices and 16 paths of length 8 each. The second Fetch environment [Plappert et al., 2018] is taken from Gymnasium-Robotics and involves controlling a 7-DoF robotic arm with a gripper attached at the end to pick up and move objects. We consider the task graph for the PickAndPlaceChoice objective that directs the gripper to move near the object, grip it, pick it up and move along one of two trajectories. The graph has 7 vertices and 2 paths of length 5. Additionally, we adapt the Rooms environment to implement a version of the drone navigation task from Example 1.
C.1.1 16-Rooms
@@ C.1 Maximizing safety during navigation and manipulation
This benchmark environment is borrowed from Jothimurugan et al. [2021]. The state and action spaces are continuous and involves controlling a point mass to navigate between rooms while avoiding obstacles. A map of the environment is shown is shown in Figure 5. Jothimurugan et al. [2021] specify objectives in the form of task graphs where each vertex corresponds to a set of subgoal states and each edge represents a reach-avoid subtask of reaching the target vertex’s subgoal states. The task graph that we consider corresponds to the specification $\phi_{5}$ from Jothimurugan et al. [2021] and is also visualized in Figure 5. It consists of 13 vertices arranged in a sequence of 4 diamonds with 4 vertices in each diamond like in Figure 1. The first diamond specifies starting from the light blue square and reaching the room marked with the dark blue square in Figure 5 using either the top path while avoiding the red obstacle or using the bottom path. Similarly, the next diamond specifies the two paths to reach the green square, followed by the pink, and brown squares. Each edge in the task graph is also associated with a real-valued reward function that evaluates to at least zero when a trajectory of the controller satisfies the reach-avoid subtask. This is function of the distance between a trajectory and the obstacles and between the trajectory and the subgoal region. So if the reward is positive then it means that the trajectory stays at least a certain distance away from the obstacles and reaches at most a certain distance to the goal.
Mapping a task graph to an agent graph is simple: to define the loss functions along each edge, we simply consider the negative of the reward function so that an upper bound on the loss corresponds to a lower bound on the reward. The agents along each edge are RL policies trained to complete the reach avoid subtasks. The policies are trained using the ARS algorithm [Mania et al., 2018] using the same recipe as described in Jothimurugan et al. [2021]. Since the training of policies are dependent on the initial state distribution, we train the policies lazily when required by the BucketedVaR algorithm by approximating the initial state distribution from samples from the previous policy.
@@
In Table 1, since the estimated VaR at all levels for the 16-Rooms benchmark is negative, it implies that with probability at least $1-\alpha$, the reward along all the edges of the path are at least the positive estimated VaR, where $\alpha$ is the risk level.
C.1.2 Fetch
The Fetch robotic arm environment [Plappert et al., 2018] from Gymnasium Robotics is a simulated robot arm environment with a gripper attached. The reach-avoid task graph is again taken from Jothimurugan et al. [2021] and has 8 vertices with two paths. It corresponds to the PickAndPlaceChoice specification and involves moving the gripper close to an object, gripping it, picking it up, and moving it along one of two trajectories. The policies are trained using TD3 [Fujimoto et al., 2018] with the hyperparameters specified by Jothimurugan et al. [2021].
C.1.3 DroneNav
@@
C.2 Minimizing resource consumption
We use the Miniworld [Chevalier-Boisvert et al., 2023] framework to implement the BoxRelay environment of which the top-view map is shown in Figure 6(a). Observations in this environment are given as first-person view RGB images (see Section C.2.1). The objective is to move the player (red triangle) to the first box (marked 1) and pick it up. The player can then choose to go to either of the two boxes in the two middle rooms (3 or 7), place the initial box, pick up the next box, and finally drop it at the location marked in the right room with a star. Furthermore, after picking up the box at 7, the player can choose to exit through one of the two exits 8 or 9. We add the constraint that the player has limited battery power but gets recharged after dropping the first box and before picking up the second box. So we would like to minimize the amount of time that either box is held by the player. Importantly, since the initial positions of the player, the boxes, and the final target are all chosen uniformly at random within each room, this is a non-trivial optimization problem.
We can model the BoxRelay task as an agent graph as follows: the graph is as shown in Figure 6(a) with the numbers being the vertices. The edges represent the agents which are RL policies trained to navigate the player to pick up and place objects at the desired locations. The loss functions evaluate to the number of time steps that a box is carried before being put back down and thus quantifies the resource consumption. Here, the RMAG objective corresponds to minimizing the maximum time that any box is carried by the player.
C.2.1 BoxRelay
(a) Top-view map with agent graph
(b) First-person view when carrying box
Figure 6: BoxRelay environment
The environment and the agent graph details are presented in Section 4. The policies are trained using the PPO [Schulman et al., 2017] CNN-Policy algorithm from StableBaselines3 package with the default hyperparameters for 500,000 iterations for each edge policy. The initial state distribution was again estimated by taking rollouts of the previous edge policy. The loss functions in the agent graph are defined as follows where the vertices are numbered as shown in Figure 6(a).
@@ C.2.1 BoxRelay
C.3 Correlated Gaussian Noise Experiment
For experiment 4, to measure the robustness of our algorithm to correlations in the loss distributions, we create a synthetic agent graph which is simply a path with a varying number of edges $p>0$. Let $C,X_{1},\dots,X_{p}\sim N(0,1)$ be i.i.d. Gaussian variables with zero mean and unit variance. Then for a given correlation level $\rho\in[0,1]$, we define the correlated edge loss variables $L_{1},\dots,L_{p}$ as
$L_{i}\coloneqq\rho C+\sqrt{1-\rho^{2}}X_{i}.$ (24)
Thus, under complete correlation $\rho=1$, all the edge loss random variables represent the same variable $C$ and under zero correlation $\rho=0$, they are independent.
C.4 Experiment computational requirements
@@
Appendix D Experimental results
Table 1: Comparison of the approximated VaR with the optimal baseline algorithm to compute VaR on $10^{4}$ samples. We prefer empirical quantiles that are close to $1-\alpha$ where $\alpha$ is the risk level. 95% Clopper-Pearson CIs are specified for the empirical quantiles computed on a fresh set of $10^{4}$ samples.
@@ Appendix D Experimental results
Table 2: Comparison of the approximated CVaR with the optimal baseline algorithm on $10^{4}$ samples. Empirical coverages are computed on a fresh set of $10^{4}$ samples.
Benchmark
Risk
level
Buckets
CVaR
estimate
Baseline
estimate
CVaR
quantile (%)
Baseline
quantile (%)
DroneNav 0.2 30 17.648 17.649 87.08 88.30
0.1 30 17.652 17.652 94.72 94.72
0.05 30 17.654 17.654 96.99 96.99
16-Rooms 0.2 100 -0.0134 -0.0139 90.82 89.75
0.1 100 -0.0077 -0.0073 95.26 94.34
0.05 100 -0.0051 -0.0045 96.55 98.70
Fetch 0.2 30 0.359 0.317 88.96 88.15
0.1 30 0.411 0.409 96.74 96.74
0.05 30 0.451 0.453 98.19 98.25
BoxRelay 0.2 50 180.54 179.86 93.57 93.49
0.1 50 210.96 210.61 96.47 96.47
0.05 50 246.30 246.27 98.24 98.24
Complete evaluation results of the BucketedVaR algorithm along with comparison with the baseline optimal algorithms to compute VaR and CVaR are given in Table 1.